# Bridging doubly heavy tetraquark mass spectrum with heavy baryons utilizing heavy antiquark-diquark symmetry

### v1

@@ III Results and discussions
The masses, quantum numbers, and configurations of the 38 predicted DHTs are systematically categorized and presented across Tables 3, 4, and 5. These encompass three distinct groups: 14 states linked to $\bar{\Lambda}_{c}$$\bar{\Lambda}_{b}$ and $\bar{\Sigma}_{c}$$\bar{\Sigma}_{b}$ baryons, 14 states associated with $\bar{\Xi}_{c}$$\bar{\Xi}_{b}$ baryons, and 10 states corresponding to $\bar{\Omega}_{c}$$\bar{\Omega}_{b}$ baryons. Each table provides comparative mass predictions from multiple theoretical approaches, revealing an overall consistency in spectral hierarchies—particularly the mass ordering $T_{QQ^{\prime}}<T_{QQ^{\prime}\bar{s}}<T_{QQ^{\prime}\bar{s}\bar{s}}$—while highlighting method-dependent variations in absolute values.
@@
Notably, our predictions derive from a framework exploiting HADS to establish rigorous connections between tetraquark spectra and experimentally measured heavy baryon spectrum. This approach contrasts with complementary methodologies: relativistic quark models employing diquark-antidiquark configurations Ebert et al. (2007) or variational methods Lü et al. (2020); effective interaction models including color-magnetic interactions Luo et al. (2017) and extended chromomagnetic model Weng et al. (2022); and non-perturbative techniques such as diffusion Monte Carlo Mutuk (2024) and MIT bag model calculations Zhang et al. (2021).
@@
Table 3: Spectrum of DHTs aroused from heavy anti-baryons $\bar{\Lambda}_{c}$($\bar{\Lambda}_{b}$) and $\bar{\Sigma}_{c}$($\bar{\Sigma}_{b}$) with HADS in this work and comparisons with different models or methods. DMC: diffusion Monte Carlo method. CMI: color-magnetic interaction model; RQM(VM): relativized quark model within the variational method; RQM(DA): relativized quark model within diquark-antidiquark picture; MIT: MIT bag model; ECM: extended chromomagnetic model.
Configuration $I(J^{P})$ This work DMC Mutuk (2024) CMI Luo et al. (2017) MIT Zhang et al. (2021) RQM(VM) Lü et al. (2020) RQM(DA) Ebert et al. (2007) ECM I Weng et al. (2022) ECM II Weng et al. (2022)
$\bar{n}\bar{n}]^{3}_{0}[cc]^{\bar{3}}_{1}]^{0}_{1}$ $0(1^{+})$ $T_{cc}(3876)$ $3892$ $4007$ $3925$ $4041$ $3935$ $3749.8$ $3868.7$
$\bar{n}\bar{n}]^{3}_{1}[cc]^{\bar{3}}_{1}]^{1}_{0}$ $1(0^{+})$ $T_{cc}(4035)$ $4062$ $4078$ $4032$ $4159$ $4056$ $3833.2$ $3969.2$
$\bar{n}\bar{n}]^{3}_{1}[cc]^{\bar{3}}_{1}]^{1}_{1}$ $1(1^{+})$ $T_{cc}(4058)$ $4104$ $4021$ $4117$ $4268$ $4079$ $3946.4$ $4053.2$
$\bar{n}\bar{n}]^{3}_{1}[cc]^{\bar{3}}_{1}]^{1}_{2}$ $1(2^{+})$ $T_{cc}(4103)$ $4207$ $4271$ $4179$ $4318$ $4118$ $4017.1$ $4123.8$
$\bar{n}\bar{n}]^{3}_{0}[bb]^{\bar{3}}_{1}]^{0}_{1}$ $0(1^{+})$ $T_{bb}(10396)$ $10,338$ $10686$ $10654$ $10550$ $10502$ $10291.6$ $10390.9$
$\bar{n}\bar{n}]^{3}_{1}[bb]^{\bar{3}}_{1}]^{1}_{0}$ $1(0^{+})$ $T_{bb}(10586)$ $10,624$ $10841$ $10834$ $10765$ $10648$ $10468.8$ $10569.3$
$\bar{n}\bar{n}]^{3}_{1}[bb]^{\bar{3}}_{1}]^{1}_{1}$ $1(1^{+})$ $T_{bb}(10593)$ $10,680$ $10875$ $10854$ $10779$ $10657$ $10485.3$ $10584.2$
$\bar{n}\bar{n}]^{3}_{1}[bb]^{\bar{3}}_{1}]^{1}_{2}$ $1(2^{+})$ $T_{bb}(10608)$ $10702$ $10897$ $10878$ $10799$ $10673$ $10507.9$ $10606.8$
$\bar{n}\bar{n}]^{3}_{0}[cb]^{\bar{3}}_{0}]^{0}_{0}$ $0(0^{+})$ $T_{cb}(7142)$ $7108$ $7256$ $7260$ $7297$ $7239$ $7003.4$ $7124.6$
$\bar{n}\bar{n}]^{3}_{0}[cb]^{\bar{3}}_{1}]^{0}_{1}$ $0(1^{+})$ $T_{cb}(7176)$ $7084$ $7321$ $7288$ $7325$ $7246$ $7046.2$ $7158.0$
$\bar{n}\bar{n}]^{3}_{1}[cb]^{\bar{3}}_{0}]^{1}_{1}$ $1(1^{+})$ $T_{cb}(7347)$ $7510$ $7473$ $7465$ $7537$ $7396$ $7211.0$ $7322.5$
$\bar{n}\bar{n}]^{3}_{1}[cb]^{\bar{3}}_{1}]^{1}_{0}$ $1(0^{+})$ $T_{cb}(7358)$ $7396$ $7457$ $7438$ $7519$ $7383$ $7189.5$ $7305.6$
$\bar{n}\bar{n}]^{3}_{1}[cb]^{\bar{3}}_{1}]^{1}_{1}$ $1(1^{+})$ $T_{cb}(7369)$ $7542$ $7548$ $7509$ $7561$ $7403$ $7211.0$ $7322.5$
$\bar{n}\bar{n}]^{3}_{1}[cb]^{\bar{3}}_{1}]^{1}_{2}$ $1(2^{+})$ $T_{cb}(7392)$ $7568$ $7582$ $7531$ $7586$ $7422$ $7293.2$ $7396.0$
@@
Table 4: Spectrum of DHTs aroused from heavy anti-baryons $\bar{\Xi}_{c}$($\bar{\Xi}_{b}$) with HADS in this work and comparisons with different models or methods. DMC: diffusion Monte Carlo method. CMI: color-magnetic interaction model; RQM(VM): relativized quark model within the variational method; RQM(DA): relativized quark model within diquark-antidiquark picture; MIT: MIT bag model; ECM: extended chromomagnetic model.
Configuration $I(J^{P})$ This work DMC Mutuk (2024) CMI Luo et al. (2017) MIT Zhang et al. (2021) RQM(VM) Lü et al. (2020) RQM(DA) Ebert et al. (2007) ECM I Weng et al. (2022) ECM II Weng et al. (2022)
$\bar{n}\bar{s}]^{3}_{0}[cc]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$ $\frac{1}{2}(1^{+})$ $T_{cc\bar{s}}(4026)$ $4056$ $4225$ $4091$ $4232$ $4143$ $3919.0$ $4051.5$
$\bar{n}\bar{s}]^{3}_{1}[cc]^{\bar{3}}_{1}]^{\frac{1}{2}}_{0}$ $\frac{1}{2}(0^{+})$ $T_{cc\bar{s}}(4127)$ $4165$ $4236$ $4165$ $4323$ $4221$ $3937.6$ $4085.7$
$\bar{n}\bar{s}]^{3}_{1}[cc]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$ $\frac{1}{2}(1^{+})$ $T_{cc\bar{s}}(4146)$ $4128$ $4363$ $4247$ $4394$ $4239$ $4073.0$ $4180.2$
$\bar{n}\bar{s}]^{3}_{1}[cc]^{\bar{3}}_{1}]^{\frac{1}{2}}_{2}$ $\frac{1}{2}(2^{+})$ $T_{cc\bar{s}}(4184)$ $4314$ $4434$ $4305$ $4440$ $4271$ $4144.3$ $4251.5$
$\bar{n}\bar{s}]^{3}_{0}[bb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$ $\frac{1}{2}(1^{+})$ $T_{bb\bar{s}}(10534)$ $10684$ $10911$ $10811$ $10734$ $10706$ $10473.1$ $10569.0$
$\bar{n}\bar{s}]^{3}_{1}[bb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{0}$ $\frac{1}{2}(0^{+})$ $T_{bb\bar{s}}(10660)$ $10721$ $10999$ $10955$ $10883$ $10802$ $10586.4$ $10684.1$
$\bar{n}\bar{s}]^{3}_{1}[bb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$ $\frac{1}{2}(1^{+})$ $T_{bb\bar{s}}(10666)$ $10736$ $11010$ $10974$ $10897$ $10809$ $10605.3$ $10700.5$
$\bar{n}\bar{s}]^{3}_{1}[bb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{2}$ $\frac{1}{2}(2^{+})$ $T_{bb\bar{s}}(10679)$ $10747$ $11060$ $10997$ $10915$ $10823$ $10628.7$ $10723.9$
$\bar{n}\bar{s}]^{3}_{0}[cb]^{\bar{3}}_{0}]^{\frac{1}{2}}_{0}$ $\frac{1}{2}(0^{+})$ $T_{cb\bar{s}}(7286)$ - $7461$ - $7483$ $7444$ $7156.5$ $7296.2$
$\bar{n}\bar{s}]^{3}_{0}[cb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$ $\frac{1}{2}(1^{+})$ $T_{cb\bar{s}}(7320)$ - $7530$ - $7514$ $7451$ $7212.8$ $7336.1$
$\bar{n}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{0}]^{\frac{1}{2}}_{1}$ $\frac{1}{2}(1^{+})$ $T_{cb\bar{s}}(7425)$ - $7634$ - $7659$ $7555$ $7323.9$ $7440.9$
$\bar{n}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{0}$ $\frac{1}{2}(0^{+})$ $T_{cb\bar{s}}(7440)$ - $7615$ - $7643$ $7540$ $7299.0$ $7421.0$
$\bar{n}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$ $\frac{1}{2}(1^{+})$ $T_{cb\bar{s}}(7449)$ - $7706$ - $7682$ $7552$ $7330.5$ $7487.8$
$\bar{n}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{2}$ $\frac{1}{2}(2^{+})$ $T_{cb\bar{s}}(7468)$ - $7718$ - $7705$ $7572$ $7415.1$ $7517.1$

### v2

@@ III Results and discussions
The masses, quantum numbers, configurations and mass formulas of the 38 predicted DHTs are systematically categorized and presented across Tables 3, 4, and 5. It should be noted that the states listed in the tables are calculated one-to-one according to configurations using the mass formula. Actually, since certain configurations share the same quantum numbers and quark components, mixing can occur between them. We can compute the correlation matrix of states with different configurations to determine the eigenmasses and eigenvectors of the mixed eigenstates, thereby obtaining the mass spectrum of DHT states considering mixing effects. The diagonal elements of the correlation matrix correspond to the masses of explicit configurations given in Tables 3, 4, and 5, while the off-diagonal elements represent the couplings between different configurations under the mass Hamiltonian (i.e., Eq.(2)). This coupling originates from the different projections of color operators and spin operators onto different configurations. Considering that we adopt HADS in this work, where the heavy diquark is treated as an integral anti-heavy quark, the configurations $QQ^{\prime}]_{1}^{3}$ and $QQ^{\prime}]_{0}^{3}$ are regarded as distinct “particles” and cannot mix. Consequently, the mixing of DHT states is simplified to a form similar to the mixing of three-quark heavy baryon configurations—this is a key distinction between the present work and other studies. By comparing the mass spectrum of mixed states with and without the HADS approximation, we can assess the impact of introducing HADS on the DHT mass spectrum.
In Table 6, we present the mass correlation matrices of different configurations under the same quantum numbers considering HADS, along with the final eigenmasses and eigenvectors. The results indicate that mixing primarily occurs for the quantum number $1/2(1^{+})$: $T_{cc}$, $T_{bb}$, and $T_{bc}$ each have two configurations that undergo mixing, with mass differences of approximately several MeV to over ten MeV compared to the unmixed cases. In the alternative scenario where we do not consider the HADS approximation and still treat DHT as a four-body system, the projection values of color operators and spin operators can be found in Tables II and III of Ref.  et al. (2020). The results for this case are listed in Table 7. Compared to the HADS-included case, $T_{cb}$ has seven additional mixed configurations, mainly due to the allowable mixing between $cb]_{1}^{\bar{3}}$ and $cb]_{0}^{\bar{3}}$ without HADS. A comparison of the results in Tables 6 and  7 shows that the introduction of HADS has an impact of approximately ten-odd MeV on the masses of mixed DHT states, consistent with the aforementioned analysis.
Table 6: The eigenmasses and eigenvectors of the DHT states with quantum numbers and mixing configurations with HADS.
$I(J^{p})$ Configuration $\langle H\rangle$ (MeV) Mass (MeV) Eigenvector
$\frac{1}{2}(1^{+})$ $\bar{n}\bar{s}]^{3}_{0}[cc]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$ $\begin{pmatrix}4026&40\\ 40&4146\end{pmatrix}$ $\begin{bmatrix}4014\\ 4158\end{bmatrix}$ $\begin{bmatrix}(-0.957,0.290))\\ (0.290,0.957)\end{bmatrix}$
$\bar{n}\bar{s}]^{3}_{1}[cc]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$
$\frac{1}{2}(1^{+})$ $\bar{n}\bar{s}]^{3}_{0}[bb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$ $\begin{pmatrix}10534&13\\ 13&10666\end{pmatrix}$ $\begin{bmatrix}10533\\ 10667\end{bmatrix}$ $\begin{bmatrix}(-0.995,0.097))\\ (0.097,0.995)\end{bmatrix}$
$\bar{n}\bar{s}]^{3}_{1}[bb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$
$\frac{1}{2}(1^{+})$ $\bar{n}\bar{s}]^{3}_{0}[cb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$ $\begin{pmatrix}7320&20\\ 20&7449\end{pmatrix}$ $\begin{bmatrix}7317\\ 7452\end{bmatrix}$ $\begin{bmatrix}(-0.989,0.150))\\ (0.150,0.989)\end{bmatrix}$
$\bar{n}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$
Table 7: The eigenmasses and eigenvectors of the DHT states with quantum numbers and mixing configurations without HADS.
$I(J^{p})$ Configuration $\langle H\rangle$ (MeV) Mass (MeV) Eigenvector
$\frac{1}{2}(1^{+})$ $\bar{n}\bar{s}]^{3}_{0}[cc]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$ $\begin{pmatrix}4026&-10\\ -10&4128\end{pmatrix}$ $\begin{bmatrix}4025\\ 4129\end{bmatrix}$ $\begin{bmatrix}(0.995,0.097))\\ (-0.097,0.995)\end{bmatrix}$
$\bar{n}\bar{s}]^{3}_{1}[cc]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$
$\frac{1}{2}(1^{+})$ $\bar{n}\bar{s}]^{3}_{0}[bb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$ $\begin{pmatrix}10534&-3\\ -3&10660\end{pmatrix}$ $\begin{bmatrix}10534\\ 10660\end{bmatrix}$ $\begin{bmatrix}(0.999,0.023))\\ (-0.023,0.999)\end{bmatrix}$
$\bar{n}\bar{s}]^{3}_{0}[bb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$
$1(1^{+})$ $\bar{n}\bar{n}]^{3}_{1}[cb]^{\bar{3}}_{0}]^{1}_{1}$ $\begin{pmatrix}7347&20\\ 20&7351\end{pmatrix}$ $\begin{bmatrix}7329\\ 7369\end{bmatrix}$ $\begin{bmatrix}(-0.741,0.671))\\ (0.671,0.741)\end{bmatrix}$
$\bar{n}\bar{n}]^{3}_{1}[cb]^{\bar{3}}_{1}]^{1}_{1}$
$\frac{1}{2}(0^{+})$ $\bar{n}\bar{s}]^{3}_{0}[cb]^{\bar{3}}_{0}]^{\frac{1}{2}}_{0}$ $\begin{pmatrix}7286&-4\\ -4&7410\end{pmatrix}$ $\begin{bmatrix}7286\\ 7410\end{bmatrix}$ $\begin{bmatrix}(0.999,0.032))\\ (-0.032,0.999)\end{bmatrix}$
$\bar{n}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{0}$
$\frac{1}{2}(1^{+})$ $\bar{n}\bar{s}]^{3}_{0}[cb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$ $\begin{pmatrix}7320&2&-7\\ 2&7425&17\\ -7&17&7434\end{pmatrix}$ $\begin{bmatrix}7319\\ 7412\\ 7447\end{bmatrix}$ $\begin{bmatrix}(0.997,-0.029,0.065))\\ (-0.063,-0.796,0.603)\\ (-0.034,0.605,0.795)\end{bmatrix}$
$\bar{n}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{0}]^{\frac{1}{2}}_{1}$
$\bar{n}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$
$(1^{+})$ $\bar{s}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{0}]_{1}$ $\begin{pmatrix}7508&14\\ 14&7522\end{pmatrix}$ $\begin{bmatrix}7499\\ 7531\end{bmatrix}$ $\begin{bmatrix}(-0.851,0.526))\\ (0.526,0.851)\end{bmatrix}$
$\bar{s}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{1}]_{1}$
@@
The masses and quantum numbers of the 38 predicted DHTs with mixing effect are categorized and presented across Tables 8, 9, and 10. These encompass three distinct groups: 14 states linked to $\bar{\Lambda}_{c}$$\bar{\Lambda}_{b}$ and $\bar{\Sigma}_{c}$$\bar{\Sigma}_{b}$ baryons, 14 states associated with $\bar{\Xi}_{c}$$\bar{\Xi}_{b}$ baryons, and 10 states corresponding to $\bar{\Omega}_{c}$$\bar{\Omega}_{b}$ baryons. Each table provides comparative mass predictions from multiple theoretical approaches, revealing an overall consistency in spectral hierarchies—particularly the mass ordering $T_{QQ^{\prime}}<T_{QQ^{\prime}\bar{s}}<T_{QQ^{\prime}\bar{s}\bar{s}}$—while highlighting method-dependent variations in absolute values. Notably, our predictions derive from a framework exploiting HADS to establish rigorous connections between tetraquark spectra and experimentally measured heavy baryon spectrum. This approach contrasts with complementary methodologies: relativistic quark models employing diquark-antidiquark configurations Ebert et al. (2007) or variational methods  et al. (2020); effective interaction models including color-magnetic interactions Luo et al. (2017) and extended chromomagnetic model Weng et al. (2022); and non-perturbative techniques such as diffusion Monte Carlo Mutuk (2024) and MIT bag model calculations Zhang et al. (2021).
@@
Table 8: Spectrum of DHTs aroused from heavy anti-baryons $\bar{\Lambda}_{c}$($\bar{\Lambda}_{b}$) and $\bar{\Sigma}_{c}$($\bar{\Sigma}_{b}$) with HADS (values in brackets are without HADS) in this work and comparisons with different models or methods. DMC: diffusion Monte Carlo method. CMI: color-magnetic interaction model; RQM(VM): relativized quark model within the variational method; RQM(DA): relativized quark model within diquark-antidiquark picture; MIT: MIT bag model; ECM: extended chromomagnetic model.
$I(J^{P})$ Configuration This work DMC Mutuk (2024) CMI Luo et al. (2017) MIT Zhang et al. (2021) RQM(VM)  et al. (2020) RQM(DA) Ebert et al. (2007) ECM I Weng et al. (2022) ECM II Weng et al. (2022)
$0(1^{+})$ $\bar{n}\bar{n}]^{3}_{0}[cc]^{\bar{3}}_{1}]^{0}_{1}$ $3876$ $3892$ $4007$ $3925$ $4041$ $3935$ $3749.8$ $3868.7$
$1(0^{+})$ $\bar{n}\bar{n}]^{3}_{1}[cc]^{\bar{3}}_{1}]^{1}_{0}$ $4035$ $4062$ $4078$ $4032$ $4159$ $4056$ $3833.2$ $3969.2$
$1(1^{+})$ $\bar{n}\bar{n}]^{3}_{1}[cc]^{\bar{3}}_{1}]^{1}_{1}$ $4058$ $4104$ $4021$ $4117$ $4268$ $4079$ $3946.4$ $4053.2$
$1(2^{+})$ $\bar{n}\bar{n}]^{3}_{1}[cc]^{\bar{3}}_{1}]^{1}_{2}$ $4103$ $4207$ $4271$ $4179$ $4318$ $4118$ $4017.1$ $4123.8$
$0(1^{+})$ $\bar{n}\bar{n}]^{3}_{0}[bb]^{\bar{3}}_{1}]^{0}_{1}$ $10396$ $10,338$ $10686$ $10654$ $10550$ $10502$ $10291.6$ $10390.9$
$1(0^{+})$ $\bar{n}\bar{n}]^{3}_{1}[bb]^{\bar{3}}_{1}]^{1}_{0}$ $10586$ $10,624$ $10841$ $10834$ $10765$ $10648$ $10468.8$ $10569.3$
$1(1^{+})$ $\bar{n}\bar{n}]^{3}_{1}[bb]^{\bar{3}}_{1}]^{1}_{1}$ $10593$ $10,680$ $10875$ $10854$ $10779$ $10657$ $10485.3$ $10584.2$
$1(2^{+})$ $\bar{n}\bar{n}]^{3}_{1}[bb]^{\bar{3}}_{1}]^{1}_{2}$ $10608$ $10702$ $10897$ $10878$ $10799$ $10673$ $10507.9$ $10606.8$
$0(0^{+})$ $\bar{n}\bar{n}]^{3}_{0}[cb]^{\bar{3}}_{0}]^{0}_{0}$ $7142$ $7108$ $7256$ $7260$ $7297$ $7239$ $7003.4$ $7124.6$
$0(1^{+})$ $\bar{n}\bar{n}]^{3}_{0}[cb]^{\bar{3}}_{1}]^{0}_{1}$ $7176$ $7084$ $7321$ $7288$ $7325$ $7246$ $7046.2$ $7158.0$
$1(0^{+})$ $\bar{n}\bar{n}]^{3}_{1}[cb]^{\bar{3}}_{1}]^{1}_{0}$ $735$ $7396$ $7457$ $7438$ $7519$ $7383$ $7189.5$ $7305.6$
$1(1^{+})$ $\bar{n}\bar{n}]^{3}_{1}[cb]^{\bar{3}}_{0}]^{1}_{1}$ $7347(7329)$ $7510$ $7473$ $7465$ $7537$ $7396$ $7211.0$ $7322.5$
$\bar{n}\bar{n}]^{3}_{1}[cb]^{\bar{3}}_{1}]^{1}_{1}$ $7369(7369)$ $7542$ $7548$ $7509$ $7561$ $7403$ $7211.0$ $7322.5$
$1(2^{+})$ $\bar{n}\bar{n}]^{3}_{1}[cb]^{\bar{3}}_{1}]^{1}_{2}$ $7392$ $7568$ $7582$ $7531$ $7586$ $7422$ $7293.2$ $7396.0$
@@
Table 9: Spectrum of DHTs aroused from heavy anti-baryons $\bar{\Xi}_{c}$($\bar{\Xi}_{b}$) with HADS (values in brackets are without HADS) in this work and comparisons with different models or methods. DMC: diffusion Monte Carlo method. CMI: color-magnetic interaction model; RQM(VM): relativized quark model within the variational method; RQM(DA): relativized quark model within diquark-antidiquark picture; MIT: MIT bag model; ECM: extended chromomagnetic model.
$I(J^{P})$ Configuration This work DMC Mutuk (2024) CMI Luo et al. (2017) MIT Zhang et al. (2021) RQM(VM)  et al. (2020) RQM(DA) Ebert et al. (2007) ECM I Weng et al. (2022) ECM II Weng et al. (2022)
$\frac{1}{2}(0^{+})$ $\bar{n}\bar{s}]^{3}_{1}[cc]^{\bar{3}}_{1}]^{\frac{1}{2}}_{0}$ $4127$ $4165$ $4236$ $4165$ $4323$ $4221$ $3937.6$ $4085.7$
$\frac{1}{2}(1^{+})$ $\bar{n}\bar{s}]^{3}_{0}[cc]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$ $4014(4025)$ $4056$ $4225$ $4091$ $4232$ $4143$ $3919.0$ $4051.5$
$\bar{n}\bar{s}]^{3}_{1}[cc]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$ $4158(4129)$ $4128$ $4363$ $4247$ $4394$ $4239$ $4073.0$ $4180.2$
$\frac{1}{2}(2^{+})$ $\bar{n}\bar{s}]^{3}_{1}[cc]^{\bar{3}}_{1}]^{\frac{1}{2}}_{2}$ $4184$ $4314$ $4434$ $4305$ $4440$ $4271$ $4144.3$ $4251.5$
$\frac{1}{2}(0^{+})$ $\bar{n}\bar{s}]^{3}_{1}[bb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{0}$ $10660$ $10721$ $10999$ $10955$ $10883$ $10802$ $10586.4$ $10684.1$
$\frac{1}{2}(1^{+})$ $\bar{n}\bar{s}]^{3}_{0}[bb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$ $10533(10534)$ $10684$ $10911$ $10811$ $10734$ $10706$ $10473.1$ $10569.0$
$\bar{n}\bar{s}]^{3}_{1}[bb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$ $10667(10660)$ $10736$ $11010$ $10974$ $10897$ $10809$ $10605.3$ $10700.5$
$\frac{1}{2}(2^{+})$ $\bar{n}\bar{s}]^{3}_{1}[bb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{2}$ $10679$ $10747$ $11060$ $10997$ $10915$ $10823$ $10628.7$ $10723.9$
$\frac{1}{2}(0^{+})$ $\bar{n}\bar{s}]^{3}_{0}[cb]^{\bar{3}}_{0}]^{\frac{1}{2}}_{0}$ $7286(7286)$ - $7461$ - $7483$ $7444$ $7156.5$ $7296.2$
$\bar{n}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{0}$ $7440(7410)$ - $7615$ - $7643$ $7540$ $7299.0$ $7421.0$
$\frac{1}{2}(1^{+})$ $\bar{n}\bar{s}]^{3}_{0}[cb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$ $7317(7319)$ - $7530$ - $7514$ $7451$ $7212.8$ $7336.1$
$\bar{n}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$ $7452(7447)$ - $7706$ - $7682$ $7552$ $7330.5$ $7487.8$
$\bar{n}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{0}]^{\frac{1}{2}}_{1}$ $7425(7412)$ - $7634$ - $7659$ $7555$ $7323.9$ $7440.9$
$\frac{1}{2}(2^{+})$ $\bar{n}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{2}$ $7468$ - $7718$ - $7705$ $7572$ $7415.1$ $7517.1$